I need to solve both of these and am able to do the first one but not the second one.
#1 is $3/lb. = c/oz
I get $3/16 = 18.5 cents per oz.
This one I need help with below
5cm/min. = ? m/week
Any ideas??
Thanks

Answer by gonzo(654)

problem 1:
-----
$3.00 per pound = ($3.00 / 16) per ounce = 18.75 cents per ounce.
-----
problem 2:
-----
5 centimeters per minute = x meters per week.
-----
first convert 1 week to minutes.
1 week = 7 days * 24 hours per day * 60 minutes per hour = 10080 minutes.
-----
since your problem states 5 centimeters per minute, and you want to know how many meters per week, you first want to get how many centimeters per week.
-----
since 1 week is 10080 minutes, and your original equation is for 1 minute, you need to multiply your equation by 10080 to find out how many centimeters in a week.
-----
5 centimeters per minute = 5 * 10080 centimeters per week = 50400 centimeters per week.
-----
now you want to convert from centimeters to meters.
100 centimeters = 1 meter, so 1 centimeter = .01 meters (1/100 of a centimeter).
to convert from centimeters to meters you have to divide the number of centimeters by 100.
-----
50400 centimeters / 100 = 504 meters.
-----
your answer is: 504 meters per week.
-----